# From Estimands to Robust Inference of Treatment Effects in Platform Trials

### v1

@@ Abstract
A platform trial is an innovative clinical trial design that uses a master protocol (i.e., one overarching protocol) to evaluate multiple treatments in an ongoing manner and can accelerate the evaluation of new treatments. However, the flexibility that marks the potential of platform trials also creates inferential challenges. Two key challenges are the precise definition of treatment effects and the robust and efficient inference on these effects. To address these challenges, we first define a clinically meaningful estimand that characterizes the treatment effect as a function of the expected outcomes under two given treatments among concurrently eligible patients. Then, we develop weighting and post-stratification methods for estimation of treatment effects with minimal assumptions. To fully leverage the efficiency potential of data from concurrently eligible patients, we also consider a model-assisted approach for baseline covariate adjustment to gain efficiency while maintaining robustness against model misspecification. We derive and compare asymptotic distributions of proposed estimators in theory and propose robust variance estimators. The proposed estimators are empirically evaluated in a simulation study and illustrated using the SIMPLIFY trial. Our methods are implemented in the R package RobinCID.
Keywords: Covariate adjustment; Eligible and concurrent controls; Estimand; Inverse probability weighting; Master protocols; Relative efficiency.
@@ 1 Introduction
The traditional paradigm of conducting a separate clinical trial for every investigational treatment has become ever more expensive and challenging (Saville and Berry,, 2016; Woodcock and LaVange,, 2017). As a result, the landscape of clinical trials has been evolving, to make clinical trials faster and more efficient while continuing to provide reliable information on safety and effectiveness (The White House Office of Science and Technology,, 2023).
The platform trial has emerged as an innovative trial design to expedite the evaluation of potentially effective treatments (Berry et al.,, 2015; Saville and Berry,, 2016; Woodcock and LaVange,, 2017; Park et al.,, 2020; Gold et al.,, 2022; Burki,, 2023; Pitre et al.,, 2023). It is a dynamic clinical trial that uses a master protocol (i.e., one overarching protocol) to evaluate multiple treatments in an ongoing manner, where new treatments can be added to the platform as they become available, and existing ones can leave the platform based on a finding of futility or efficacy. Such a trial structure not only streamlines operations but also has the potential to improve statistical efficiency by facilitating the sharing of data between different treatment evaluations (Mehta et al.,, 2023). The effectiveness of platform trials in addressing urgent medical challenges quickly is evidenced by the hallmark RECOVERY platform trial, which was designed to rapidly evaluate COVID-19 therapies during the pandemic (Normand,, 2021; Buenconsejo et al.,, 2023). Platform trials are also attractive in the exploratory phase to quickly screen multiple treatments for diseases (Wang and Yee,, 2019) and in rare diseases to enhance the chance of identifying effective treatments (Dhaenens et al.,, 2024). Two examples of platform trials are presented in Section 2 to provide detailed illustrations.
Although platform trials offer several advantages over traditional trials, they also create challenges in statistical inference on treatment effects, as highlighted in a recent Food and Drug Administration draft guidance (FDA,, 2023). One primary challenge is how to define a meaningful treatment effect in a platform trial setting, as multiple treatments may target different biomarker-defined subgroups and enter the trial at different times — a complexity that does not exist in traditional trials. In platform trials, when comparing a control and a treatment of interest, many researchers (Lee and Wason,, 2020; Dodd et al.,, 2021) and FDA, (2023) have recommended using eligible and concurrent controls, which refer to the patients who are eligible to receive the treatment and are concurrently randomized to the control arm during the same enrollment period as the other patients being assigned to the treatment arm. The use of eligible and concurrent controls, or concurrently eligible patients, preserves the integrity of randomized comparisons and inference on treatment effects. In contrast, patients who are either ineligible or nonconcurrent may not be comparable to the treated group due to measured and unmeasured differences (FDA,, 2023). In Section 3, we explicitly present, for the first time in the literature, how to construct a clinically meaningful estimand with clear specificity on the population of concurrently eligible patients, which aligns with the ICH E9 , 2019 (R1) estimand framework.
The next challenge is on robust estimation of the defined treatment effects for concurrently eligible patients. Given the adaptive nature of platform trials, where treatments can be added and removed, the available treatment options vary over time. This dynamic, coupled with different eligibility criteria for each treatment, means that the probabilities of treatment assignments can vary over time and patient characteristics — another complexity absent in traditional trials. Thus, naively comparing outcome means between two treatment arms may yield biased results, an example of Simpson’s paradox (Collignon et al.,, 2020; Park et al.,, 2020; Dodd et al.,, 2021; Lee et al.,, 2021; Meyer et al.,, 2021).
The literature on statistical methods for platform trials is expanding, though it primarily focuses on model-based techniques for incorporating nonconcurrent controls, which can be sensitive to model misspecification (Saville et al.,, 2022; Bofill Roig et al.,, 2022; Elm et al.,, 2012; Wang et al.,, 2023). Some research has been conducted on statistical methods for concurrently eligible patients. Ventz et al., (2018) considered several adaptive randomization procedures and tests for the hypothesis of no treatment effect. Marschner and Schou, (2022) proposed stratification using enrollment windows, without taking into account other factors related to treatment and randomization. Huang et al., (2023) developed inference methods for time-to-event outcomes under a strong constancy assumption that conditional on a discrete baseline covariate, the efficacy of treatments is invariant across enrollment windows. However, there remains a lack of rigorous investigation on how to make robust inference on treatment effects in platform trials under the same minimal statistical assumptions as those in traditional clinical trials.
To fill this gap, in Section 4, we develop methods for estimating treatment effects for concurrently eligible patients using inverse probability weighting or post-stratification to account for varying treatment assignment probabilities. Additionally, we consider model-assisted approaches for covariate adjustment to gain efficiency while maintaining robustness against model misspecification. In Section 5, we theoretically establish the asymptotic distributions of proposed estimators and compare their asymptotic relative efficiency. We provide robust variance estimators in Section 6. Simulation results and an application of our methods to the SIMPLIFY trial (Mayer-Hamblett et al.,, 2023) are in Section 7. Our methods are implemented in the R package RobinCID.
2 Examples of Platform Trials
2.1 A stylistic example
First, we present a stylistic example of a platform trial in Figure 1 to provide a concrete illustration of its key features. This trial comprises five sub-studies, each containing a randomized controlled comparison of several treatments. Initiating or stopping a sub-study is governed by a master protocol. Additionally, each sub-study has an intervention-specific appendix outlining protocol elements specific to that sub-study. Patients are randomized first to eligible sub-studies and subsequently to treatment arms within a sub-study, which gives explicit treatment allocation probabilities for every subject, sufficient for the identification of treatment effects and statistical analysis. Real trials resembling this structure include Lung-MAP for lung cancer (Herbst et al.,, 2015; Redman et al.,, 2020) and an ongoing trial for chronic pain (ClinicalTrials.gov, NCT05986292).
Figure 1: A stylistic example of a platform trial. The x-axis represents calendar time discretized into windows with different potential enrollment in components of the platform trial. The figure illustrates both different treatment comparisons and different patient subgroups evolving over time.
In Figure 1, each box represents a sub-study along with its status, and each striped block within a box represents the enrollment period for a treatment evaluated in that sub-study. The horizontal axis denotes calendar time. Enrollment windows are defined by changes in the available treatments for enrollment. For instance, the initiation of enrollment for sub-study 3 marks the start of enrollment window 2, within which treatment 4 is compared to treatment 1 as the control among patients with disease subtype A. Likewise, the conclusion of the enrollment period for sub-study 2 defines the start of enrollment period 3. Subsequently, the start of treatment 3 as the new control in place of treatment 1 among subtype A defines the start of enrollment window 4.
The following is a narrative of Figure 1 according to the order of enrollment windows.
Enrollment Window 1. In this example, there are two initial sub-studies. Patients with any disease subtype is eligible to sub-study 1, while sub-study 2 is restricted to patients with subtype A. During enrollment window 1, patients with subtype A are randomly assigned to sub-study 1 or 2, and then randomized to treatment 1 or 2 if they are in sub-study 1 and to treatment 1 or 3 if they are in sub-study 2. Patients with disease subtypes other than A are enrolled in sub-study 1 and then randomized to treatment 1 or 2.
Enrollment Window 2. Sub-study 3 is added to the platform trial when a new intervention treatment 4 becomes available for disease subtype A. Following this addition, a patient with subtype A is randomly assigned to either sub-study 1, 2, or 3, and subsequently randomized to the treatments being assessed within the respective sub-study.
Enrollment Window 3. After sub-study 2 reaches full enrollment, it closes its enrollment. Later during this period, treatment 3 meets the pre-specified success criteria.
Enrollment Window 4. Since treatment 3 has been demonstrated to be superior to treatment 1 for subtype A, it replaces treatment 1 as the control for subtype A. Consequently, treatment 3 replaces treatment 1 in sub-study 3. Moreover, sub-study 1 excludes patients with subtype A, and sub-study 4 is introduced to continue evaluating treatment 2 in comparison to the new control treatment 3 for patients with subtype A.
Enrollment Window 5. Sub-study 5 is added upon the introduction of a new treatment 5 for patients positive for biomarker B. In sub-study 5, treatment 5 is evaluated against treatment 1 for patients without subtype A, or to treatment 3 for patients with subtype A.
Enrollment Window 6. Sub-study 4 has finished its planned enrollment and is no longer accepting new patients.
Enrollment Window 7. A new treatment, treatment 6, has been introduced in sub-study 5 for patients positive for biomarker B. During this enrollment window, treatment 2 in sub-study 4 fails to meet the pre-specified success criteria.
As the platform trial continues, additional sub-studies or treatments may be added as new interventions emerge for one or more targeted groups of patients, while existing sub-studies may be concluded due to either meeting success criteria or demonstrating futility.
As exemplified in Figure 1, platform trials can be very flexible: adding new sub-studies and/or treatments with different eligibility, dropping treatments or sub-studies, and changing the control in the middle of the trial. Moreover, compared to a traditional clinical trial, it enables potential pooling of data across sub-studies. Specifically, in traditional clinical trials, patients self-select into studies and are randomized into treatment arms, with treatment comparisons conducted solely within each study. In contrast, platform trials enroll patients into a master protocol, randomizing them into eligible sub-studies and subsequently treatment arms, and consequently, patients can be pooled across sub-studies. For example, when comparing treatments 1 and 2, in a platform trial like Figure 1, patients from not only sub-study 1 but also other sub-studies receiving treatment 1 can be included in the analysis, whereas we analyze data within sub-study 1 in a traditional trial. See the simulation results and discussions in Section 7.1 for the efficiency gain.
2.2 A real example: the SIMPLIFY trial
@@
Our second example is a real platform design aliased SIMPLIFY for people with cystic fibrosis (Mayer-Hamblett et al.,, 2023). It contains two concurrently run multi-center randomized controlled studies at 80 participating clinics across the United States in the Cystic Fibrosis Therapeutics Development Network. The primary objective for each study is to determine whether discontinuing either hypertonic saline (HS) or dornase alfa (DA), respectively, is non-inferior to continuing therapy in individuals using the highly effective modulator, elexacaftor plus tezacaftor plus ivacaftor (ETI). The outcome is the 6-week mean absolute change in percent-predicted forced expiratory volume in one second (ppFEV1). Figure 2 depicts the SIMPLIFY study design.
@@
Figure 2: SIMPLIFY study design. All enrolled individuals have initiated ETI for at least 90 days and belong to one of three groups: those receiving HS only, those receiving DA only, and those receiving both therapies. Individuals on a single therapy enter the corresponding study, while those on both are randomized to the HS or DA study, first with a 1:1 probability and then with a 3:1 probability during subsequent enrollment. All individuals in HS study (or DA study) are randomized 1:1 to continue or discontinue HS (or DA).
3 Setup and Estimand
3.1 Randomization procedures in platform trials

### v4

@@ Abstract
A platform trial is an innovative clinical trial design that uses a master protocol (i.e., one overarching protocol) to evaluate multiple treatments, where patients are often assigned to different subsets of treatment arms based on individual characteristics, enrollment timing, and treatment availability. While offering increased flexibility, this constrained and non-uniform treatment assignment poses inferential challenges, with two fundamental ones being the precise definition of treatment effects and robust, efficient inference on these effects. Such challenges arise primarily because some commonly used analysis approaches may target estimands defined on populations inadvertently depending on randomization ratios or trial operation format, thereby undermining interpretability. This article, for the first time, presents a formal framework for constructing a clinically meaningful estimand with precise specification of the population of interest. Specifically, the proposed entire concurrently eligible (ECE) population not only preserves the integrity of randomized comparisons but also remains invariant to both the randomization ratio and trial operation format. Then, we develop weighting and post-stratification methods to estimate treatment effects under the same minimal assumptions used in traditional randomized trials. We also consider model-assisted covariate adjustment to fully unlock the efficiency potential of platform trials while maintaining robustness against model misspecification. For all proposed estimators, we derive asymptotic distributions and propose robust variance estimators and compare them in theory and through simulations. The SIMPLIFY trial, a master protocol assessing continuation versus discontinuation of two common therapies in cystic fibrosis, is utilized to further highlight the practical significance of this research. All analyses are conducted using the R package RobinCID.
Keywords: Concurrently eligible individuals; Covariate adjustment; Estimand; Inverse probability weighting; Master protocols; Relative efficiency.
@@ 1.1 Background
The traditional paradigm of conducting a separate clinical trial for every investigational treatment has become ever more expensive and challenging (Woodcock and LaVange,, 2017). As a result, the landscape of clinical trials has been evolving, to make clinical trials faster and more efficient while continuing to provide reliable information on safety and effectiveness (The White House Office of Science and Technology,, 2023).
Master protocol trials have emerged as innovative designs to accelerate the evaluation of potentially effective treatments (Woodcock and LaVange,, 2017). These trials evaluate one or more treatments across one or more diseases within an overarching study structure. Three major types of master protocol designs are umbrella trials, platform trials, and basket trials. Umbrella trials evaluate multiple therapies for a single disease (Rashdan and Gerber,, 2016; Renfro and Sargent,, 2017), basket trials evaluate a medical product for multiple diseases or disease subtypes (Redig and Jänne,, 2015; Hobbs et al.,, 2022), while platform trials offer the added flexibility of adding or removing treatment arms over time (Berry et al.,, 2015; Saville and Berry,, 2016; Park et al.,, 2020; Gold et al.,, 2022; Burki,, 2023; Pitre et al.,, 2023). These three types are not mutually exclusive, and a master protocol trial can incorporate features of more than one type. For example, an umbrella-platform hybrid for lung cancer might test different treatments across biomarker-defined subtypes (umbrella) while allowing treatments to enter or exit over time (platform). Similarly, a basket-platform hybrid might study a single targeted therapy across multiple cancers (basket), with new cancer types or biomarkers added adaptively over time (platform). Therefore, umbrella and basket trials can be viewed as special cases of platform trials, and, in this paper, we use the terms platform trial and master protocol interchangeably.
Master protocols originated in oncology and remain widely used in this field, where many new biomarker-driven therapies require efficient evaluation (Park et al.,, 2019, 2020; Ouma et al.,, 2022). By utilizing a common screening platform to efficiently screen for biomarkers and identify all treatments for which a patient is eligible, these trial designs significantly improve operational efficiency, accelerate patient accrual, and enable the study of rare tumor subtypes within one trial (Woodcock and LaVange,, 2017). In addition to streamlining operations, these designs also have the potential to improve statistical efficiency by facilitating data sharing across different treatment evaluations. Notable examples include the BATTLE umbrella trials (Kim et al.,, 2011; Papadimitrakopoulou et al.,, 2016), the Lung-MAP platform trial (Herbst et al.,, 2015), and the I-SPY 2 platform trial (Barker et al.,, 2009). The utility of master protocol trials is much broader. They are gaining prominence in other diseases areas, including infectious diseases (Normand,, 2021), neurodegenerative diseases such as amyotrophic lateral sclerosis (Quintana et al.,, 2023) and Alzheimer’s disease (Bateman et al.,, 2017), as well as inflammatory bowel disease (Honap et al.,, 2024). Section 2 presents two trials to provide concrete illustrations.
1.2 Two fundamental gaps
Compared to traditional parallel-group trials, master protocol trials often share a key feature: constrained and non-uniform treatment assignment. Constrained means some participants have zero probability of receiving certain treatments, while non-uniform refers to varying non-zero assignment probabilities based on participant characteristics, enrollment time, or treatment availability. This feature marks a key distinction between master protocol and traditional trial designs, and poses significant challenges for the interpretability of and statistical inference for the treatment effect, some of which are highlighted in a recent draft guidance from the U.S. Food and Drug Administration (FDA, 2023b, ). Although master protocols are increasingly conducted, two fundamental questions arising from their constrained and non-uniform treatment assignment remain unresolved.
1. How to define a clinically meaningful estimand?
An estimand is “a precise description of the treatment effect reflecting the clinical question posed by a given clinical trials objective” and is of ultimate importance (ICH E9 , 2019, R1). In master protocol trials, treatments often have different eligibility criteria and can be added or removed over time, leading to changes in the set of available treatment options throughout the trial. Consequently, the set of treatments to which an individual can be randomized and the corresponding assignment probabilities can vary over time and across individual characteristics. This leads to constrained and non-uniform treatment assignments. As a result, defining an estimand becomes more challenging because the treatment effect must be interpreted in the context of restricted eligibility, variable assignment probabilities, and evolving trial structure over time. In particular, among the five estimand attributes (ICH E9 , 2019 (R1)), the target population of interest requires special attention. Many researchers (Lee and Wason,, 2020; Dodd et al.,, 2021) and FDA, 2023b have recommended using a population of individuals who are concurrently randomized and meet the eligibility criteria, as this preserves the integrity of randomized comparisons and avoids systematic differences between treatment groups due to measured and unmeasured factors. However, this general recommendation does not uniquely define the population, leading to ambiguity. For instance, in Cases I-II in Section 3.2, a seemingly reasonable population (and thus the estimand) may depend arbitrarily on randomization probabilities, even though it technically follows the general recommendation in FDA, 2023b .
To address this question, Section 3.2 introduces a population for comparing two treatments (typically with one as a control) that not only follows the general recommendations in FDA, 2023b but also respects the eligibility criteria while avoiding arbitrary dependencies on randomization probabilities. We refer to this as the entire concurrently eligible (ECE) population. This definition also applies to traditional parallel-group trials, to which we draw simple connections using Case I in Section 3.2 to explain why this population is sensible. The corresponding estimand for the ECE population is defined in Section 3.3. As in traditional trials, the target population in a master protocol trial may sometimes be a specific subgroup or a re-weighted population. In such cases, the ECE population can serve as the reference and may be restricted to the relevant subgroup or re-weighted to resemble the target population. Our proposal is the first in the literature to offer a systematic framework for defining clinically meaningful estimands with clear specification of the population of interest. Given the central importance of the estimand, our contribution enhances the fundamental understanding of master protocol trials and lays the groundwork for future statistical research—both frequentist and Bayesian—on topics such as interim analysis and the use of nonconcurrent controls.
2. How to provide robust estimation and inference regarding the defined treatment effect?
Although randomization is preserved, the commonly used estimator in traditional trials—comparing outcome means between treatment arms—can easily lead to biased estimates in master protocol trials, where treatment assignment is often constrained and non-uniform, potentially giving rise to issues such as Simpson’s paradox (Collignon et al.,, 2020; Dodd et al.,, 2021; Lee et al.,, 2021; Meyer et al.,, 2021). Consequently, estimating treatment effects in master protocol trials is not straightforward.
A promising first step is to adapt inverse probability weighting or post-stratification to account for the constrained and non-uniform treatment assignment. Additionally, model-assisted covariate adjustment can offer another means to further unlock the efficiency potential of master protocol trials, while maintaining robustness against model misspecification. To guide practitioners in selecting among these methods, we derive the asymptotic variance and relative efficiency of these estimation methods. These results are novel and, to our knowledge, have not previously appeared in the literature.
1.3 Prior work
The increasing popularity of platform trials has led to a surge of research addressing their statistical challenges in both design and analysis. Many studies have focused on adaptive designs, which aim to efficiently screen candidate treatments and adaptively randomize individuals to the most promising treatments (Yuan et al.,, 2016; Kaizer et al.,, 2018; Ventz et al.,, 2018). Additionally, Bofill Roig et al., (2024) examined the optimal allocation ratios in these trials. A key concern in the analysis of platform trials is the potential presence of time trends, which can introduce bias when using nonconcurrent controls (Bofill Roig et al.,, 2023). To address this, various strategies have been proposed to account for time trends in the analysis (Elm et al.,, 2012; Saville et al.,, 2022; Bofill Roig et al.,, 2022; Jiang et al.,, 2023; Wang et al.,, 2023; Guo et al.,, 2024). However, all of these statistical methods have predominantly relied on correct modeling and are sensitive to model misspecification.
The platform trial literature on non-model-based approaches for concurrently eligible individuals remains limited. Marschner and Schou, (2022) proposed stratification by enrollment windows. Huang et al., (2023) developed inference methods for time-to-event outcomes under a strong constancy assumption that, conditional on a discrete baseline covariate, the efficacy of treatments is invariant across enrollment windows. More recently, Santacatterina et al., (2025) defined an estimand based on the concurrent period and established identification and estimation results using only concurrent controls. They also proposed strategies to incorporate nonconcurrent controls under a strong conditional exchangeability assumption between concurrent and nonconcurrent controls, given observed covariates. However, none of these studies consider the more general setting in which treatments have distinct eligibility criteria, nor do they rigorously investigate how to define a clinically meaningful estimand within a general framework that is independent of specific study designs or modeling assumptions.
1.4 Outline of the article
Section 2 presents the SIMPLIFY trial and a stylized trial used to emphasize key issues. Section 3 details the ECE population and corresponding estimand. In Section 4, we develop robust estimation methods, without or with covariate adjustment. In Section 5, we establish the asymptotic distributions and relative efficiency of the proposed estimators. We highlight the practical significance through the SIMPLIFY trial in Section 6, and we conclude with a summary and recommendations in Section 7.
2 Real and Stylistic Trials
2.1 The SIMPLIFY trial
We begin by presenting the real master protocol trial that motivates this work: the SIMPLIFY trial (Mayer-Hamblett et al.,, 2023). Although it resembles an umbrella trial, where multiple treatments are evaluated concurrently and have different eligibility criteria, it also presents a key feature of platform trials: varying randomization ratio over time. Despite its seemingly simple design diagram, the trial involves both constrained and non-uniform treatment assignments. Its clarity makes it ideal for highlighting the two fundamental gaps, as it facilitates understanding of the underlying challenges.
@@ 2.1 The SIMPLIFY trial
Cystic fibrosis (CF) is caused by mutations in the CF transmembrane conductance regulator (CFTR) gene. CFTR modulator therapies are a class of small-molecule drugs designed to address the underlying cause of CF, leading to transformative improvements in patient outcomes. A landmark example is the triple-combination therapy elexacaftor/tezacaftor/ivacaftor (ETI), approved for individuals with at least one F508del allele, a mutation present in over 85% of people with CF. Given the potential overlap in therapeutic effects between ETI and pre-existing treatments such as hypertonic saline (HS) and dornase alfa (DA) and the need to reduce treatment burden, it is important to evaluate the impact of discontinuing these pre-existing therapies in individuals taking ETI. The SIMPLIFY trial was conducted within this context. It includes two concurrently run multi-center randomized controlled studies at 80 participating clinics across the United States in the CF Therapeutics Development Network. The primary objective for each study is to determine whether discontinuing either HS or DA, respectively, is non-inferior to continuing therapy in individuals using the highly effective modulator ETI. The outcome is the 6-week mean absolute change in percent-predicted forced expiratory volume in one second (ppFEV1).
@@
Figure 1: SIMPLIFY design schema. Individuals on a single therapy enter into the corresponding study with probability 1, while those on both are randomized to the HS or DA study, with ratio of 1:1 during enrollment window (EW) 1, then 3:1 during EW 2. Within HS (or DA) study, individuals are randomized 1:1 to continue or discontinue HS (or DA).
Table 1: Assignment probabilities in the SIMPLIFY trial.
Baseline variables EW Studies Treatments
HS DA EW HS study DA study 1 2 3
1 1 EW1 0.50 0.50 0.50 0.25 0.25
1 1 EW2 0.75 0.25 0.50 0.375 0.125
1 0 any 1 0 0.50 0.50 0
0 1 any 0 1 0.50 0 0.50
Figure 1 depicts the SIMPLIFY study design. All enrolled individuals have initiated ETI for at least 90 days and belong to one of three groups based on medication use when entering the trial: those on HS only, DA only, or both therapies. Individuals on a single therapy (HS only or DA only) are assigned to the corresponding study with probability 1. Those on both therapies are randomized to the HS or DA study, with 1:1 ratio during initial enrollment window 1 (before January 14, 2022), and 3:1 ratio during enrollment window 2 (after January 14, 2022) to assign more participants to the HS study, as the DA study was enrolling more quickly. This swift shift in the randomization ratio exemplifies how the flexibility of master protocol trials can accelerate clinical trials. Within each study (HS or DA), participants were then randomized 1:1 to either continue or discontinue their respective therapy. Table 1 summarizes assignment probabilities of the SIMPLIFY trial. It illustrates both constrained and non-uniform treatment assignment: participants receiving only HS or DA have zero probability of being assigned to certain treatment arms, and the non-zero probabilities of receiving any given pair of treatments may vary based on baseline characteristics and enrollment windows. Detailed Analysis is carried out in Section 6.
2.2 A stylistic trial
To motivate a general statistical framework for master protocol trials, we introduce a stylistic trial in two operation formats, both can be unified under our proposed framework. This unification highlights an important point: while trial design may influence clinical operations, it should not alter the clinical questions being addressed. Figure 2(a) presents the stylistic trial in a sub-study format, where treatments are evaluated in separate sub-studies. Figure 2(b) shows a multi-arm format, where each treatment is studied in a separate arm without dividing the trial into sub-studies.
The master protocol trial (in sub-study format) in Figure 2(a) comprises four sub-studies, each containing a randomized controlled comparison of two treatments. Initiating or stopping a sub-study is governed by a master protocol. Additionally, each sub-study has an intervention-specific appendix outlining protocol elements specific to that sub-study. Individuals are randomized first to eligible sub-studies and subsequently to treatment arms within a sub-study, which gives explicit treatment allocation probabilities for every subject, sufficient for the identification of treatment effects and statistical analysis. Real trials resembling this structure include Lung-MAP for lung cancer (Herbst et al.,, 2015; Redman et al.,, 2020) and an ongoing trial for chronic pain (ClinicalTrials.gov, NCT05986292). A narrative of Figure 2(a) is included in Section S1 of the Supplement.
Figure 2: A stylistic master protocol trial in two operating formats: (a) sub-study format, (b) multi-arm format. Each box in (a) represents a sub-study, along with its eligibility criteria and status. Each striped block within a box in (a) or standalone in (b) represents the enrollment period for a treatment evaluated. The horizontal axis shows calendar time, divided into discrete windows based on changes in available treatments for enrollment.
In Figure 2(b), we reframe the same master protocol trial into a multi-arm format. In enrollment window 1, there are three initial treatment arms: investigational treatments 2 and 3 and treatment 1 serving as the common control. In this window, individuals with subtype A can be randomized to any of the three arms, while those with subtypes other than A can only be randomized to treatments 1 or 2. Descriptions for subsequent enrollment windows are omitted due to their similarity with Figure 2(a). Examples of real trials that follow this structure include the STAMPEDE trial for prostate cancer (Sydes et al.,, 2012) and the RECOVERY trial for COVID-19 therapies (Normand,, 2021).
As exemplified in Figure 2, master protocol trials can be very flexible: adding new sub-studies and/or treatments with new interventions and different eligibility, dropping treatments or sub-studies, and changing the control in the middle of the trial.
3 Estimand
3.1 Randomization procedure